Operation Summary for March 24, 2010 - March 26, 2010 |
Wednesday March 24 The day shift began with Operations monitoring the Tevatron (TeV) store 7704 (109.33E30) and the Recycler stash (241.8E10), with the Antiproton Source stacking (3.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, one antiproton transfer was lost when a power supply (M:V105) did not ramp properly. During the evening shift, Operations terminated store 7704 at 5:54 PM. RF personnel repaired a TeV RF station (TRF8) and brought it back online after the store was terminated. Operations disabled the MTest beam at 6:05 PM. Operations put the TeV into shot setup and established store 7705 at 7:01 PM with an average initial luminosity (AIL) of 376.4E30. A MI RF station (MIRF18) tripped off several times and then wouldn’t reset. Operators bypassed the station and brought MIRF15 on in its place. Operations reported some downtime due to Linac trips (LRF5 and the Klystron Vernier). Thursday March 25 The midnight shift began with Operations monitoring store 7705 (163.68E30) and the stash (140.4E10), with Pbar stacking (8.8E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 5:51 AM. Operations reported some downtime when the I- Source SRM failed, which inhibited beam out of the source. The source recovered, but its output was down by about 50%. Operations consulted the source expert and he came in to investigate. He had recovered the source with almost all the lost intensity by 6:30 AM. During the day shift, Operations terminated store 7705 at 1:55 PM and put the TeV into shot setup. Operations established store 7706 at 3:18 PM with an AIL of 348E30. Cryo system personnel worked on the sectors C4 wet engine and controls personnel worked on the TeV C1 BPM code. MI personnel restored the MI-10 LCW system to its semi-normal mode (LCW okay, but no backup pond water pump). During the evening shift, Operations reported that problems with TeV vacuum and cryo system eventually resulted in the termination of store 7706 at 6:01 PM. Operators prepared the TeV and D0 for access. Mechanical technicians changed out the LCW filters at A4 and F3 due to differential pressure alarms. Operations disabled the MTest beam at 6 PM. Operations reported some stacking downtime due to problems with DRF1-7, which tripped due to a blown fuse on the fan filament control chassis. Friday March 26 The midnight shift began with the TeV turn on in progress, with Operations monitoring the stash (225.3E10), with Pbar stacking (19.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TeV through a dry squeeze at 1 AM, followed by a wet squeeze, and orbit corrections. Operations put the TeV into shot setup at 3:38 AM. The shot setup was short-lived. A Booster kicker (MKS09) failed. Operations called in kicker personnel and they reterminated a faulted cable. Operations put the TeV into shot setup at 6:51 AM and established store 7711 at 7:51 AM with an AIL of 252.9E30. Another Booster kicker (MKS11) tripped off just before the end of the shift; kicker personnel investigated. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/26/2010 email Fermilab |
| Security, Privacy, Legal | |